OpenStudy (mackenzie2013):

What is the surface area of a rectangular prism with the following dimensions? length: 8 cm width: 13 cm height: 10.5 cm

OpenStudy (anonymous):

indeed its 649..

OpenStudy (mackenzie2013):

If all the dimensions of the prism are doubled, what will be its new surface area? I dont understand this.. Someone explain?

OpenStudy (anonymous):

You've to use this formula:\[S=2*(Height\times Length+Height\times Width+Width\times Length)\\=2*(H\times L+H\times W+W\times L)\]\[S=2*(8\times13+8\times10.5+13\times10.5)\]\[\boxed{S=649}\]

OpenStudy (anonymous):

@Mackenzie2013 this might help.. https://www.google.com/#q=surface+area+of+a+rectangular+prism

OpenStudy (anonymous):

@Mackenzie2013 got it??

OpenStudy (anonymous):

If all the dimensions are doubled, \[S'=2*(2H\times2W+2H\times 2L+2W\times 2L)\]\[S'=2*(2*8\times2*13+2*8\times 2*10.5+2*10.5\times 2*13)\]The answer will be:\[\boxed{S'=2596}\]

OpenStudy (anonymous):

\[cm^2\]

OpenStudy (mackenzie2013):

Thank you guys!